Calculate the number of moles and the mass of Kl in 325 male of 8.23×10^-5 M Kl
1
Expert's answer
2019-11-01T05:45:05-0400

325 mL=0.325 L

n(KI)=0.325 L*8.23*10-5 mole/L=2.67*10-5 moles

m(KI)=2.67*10-5 mole*166 g/mole=4.43*10-3 g